# TR-LLM: Integrating Trajectory Data for Scene-Aware LLM-Based Human Action Prediction

### v1

@@ Abstract
Visit our project page for details (https://sites.google.com/view/trllm/%E3%83%9B%E3%83%BC%E3%83%A0).
@@ III-C2 Trajectory-based prediction
We utilize past trajectory data and scene map information to constrain the target area of a person. Since the target area prediction is heavily influenced by factors such as room layout, the person’s location, and speed, we leverage the LocoVR dataset to learn these intricate relationships. The LocoVR dataset consists of over 7,000 human trajectories captured in 131 indoor environments, with each trajectory segment spanning from a randomly assigned start to a goal location within the room. We employed a simple U-Net model to learn how to specify the possible target area based on the past trajectory and scene map (see Appendix-B2). Once the model predicts the probabilistic distribution of the target area, we calculate the overlap between this predicted distribution and the spatial location of objects in the scene to estimate the probability of each object being the target.
@@ IV-A Evaluation data
Furthermore, we manually designed daily action scenarios, detailing factors such as day/time, persona, current location, target object, subsequent actions, conversations, and more. These scenarios were transformed into text prompts for input to the LLM. We generated a trajectory from the current location to the target object by leveraging a model trained on the LocoVR dataset, which incorporates the scene map to guide the path (see Appendix-C2 for details).

### v2

@@ Abstract
Project page:
https://sites.google.com/view/trllm?usp=sharing
Github repo:
https://github.com/kojirotakeyama/TR-LLM/blob/main/readme.md
@@ III-C2 Trajectory-based prediction
We utilize past trajectory data and scene map information to constrain the target area of a person. Since the target area prediction is heavily influenced by factors such as room layout, the person’s location, and speed, we leverage the LocoVR dataset[32] to learn these intricate relationships. The LocoVR dataset consists of over 7,000 human trajectories captured in 131 indoor environments, with each trajectory segment spanning from a randomly assigned start to a goal location within the room. We employed a simple U-Net model to learn how to specify the possible target area based on the past trajectory and scene map (see Appendix-B2). Once the model predicts the probabilistic distribution of the target area, we calculate the overlap between this predicted distribution and the spatial location of objects in the scene to estimate the probability of each object being the target.
@@ IV-A Evaluation data
Furthermore, we manually designed daily action scenarios, detailing factors such as day/time, persona, current location, target object, subsequent actions, conversations, and more. These scenarios were transformed into text prompts for input to the LLM. We generated a trajectory from the current location to the target object by leveraging a model trained on the LocoVR dataset[32], which incorporates the scene map to guide the path (see Appendix-C2 for details).